If 15.00 mL of bleach (contains HOCl) requires 45.2 mL of 0.200 M KOH to reach the equivalence point, calculate the number of grams of HOCl contained in a 1.0 L bottle of bleach. Show all steps.
1
Expert's answer
2020-03-31T13:43:54-0400
  • At the equivalence point equality is observed for reacting substances: "C(1)\u22c5V(1)=C(2)\u22c5V(2)" , in particular for this question: "C(HClO)\u22c5V(HClO)=C(KOH)\u22c5V(KOH)". On the basis of this equality, it`s necessary to find C(HClO):

C(HClO) = "(C(KOH)\u22c5V(KOH))\/V(HClO)" = (0.2(M) ⋅ 0.05(L))/0.015(L) = 0.67 M

  • The next step is to find the amount of substance of HClO with concentration 0.67M in 1 L:

"\nu(HClO) = C(HClO) ⋅ V(HClO)" = 0.67(M) ⋅ 1(L) = 0.67 mol

  • Now, we can find the mass of HClO in 1 L:

"m(HClO) = \nu ⋅ M" = 0.67(mol) ⋅ 52.46(g/mol) = 35.15 g

Answer: m(HClO)=35.15 g.